# An analysis of debris-induced nonlinear phenomena in a positive ion-negative ion (PINI) plasma

### v1

@@ V Debris velocity and nonlinearity
One very interesting phenomenon is the changing property of the nonlinearity with increasing debris velocity which is similar to what is observed experimentally with respect to beam velocity in a PINI plasma [4]. What we observe is that as debris velocity increases, the wave amplitude steadily increases until reaching a maximum and then decreases, which is seen from the steepening in the leading edge (precursor region), which then gradually changes to a DSW before becoming a train of solitons and then a single KdV soliton. While the steepening in the leading edge indicates a dominant positive nonlinearity over dispersion at low debris velocity, as debris velocity increases, nonlinearity decreases and gets balanced by dispersion which creates a KdV soliton. In Fig.6, we show this effect as obtained from the fKdV solution (left column) and from mFCT hydrodynamic simulation (right column). In this figure, we have indicated the changing nonlinearity with an arrow with the keywords ‘steepening’, ‘DSW’, ‘train’, and ‘KdV’, respectively to indicate the phenomenon of steeping in the leading edge, formation of DSW, transformation to a train of solitons, and finally a simple KdV soliton as debris velocity $v_{d}$ increases from $0.8$ to $1.5$ for fKdV solutions and from $0.3$ to $1.8$ for mFCT solutions. All the panels are drawn at the rest frame of the debris, which is moving toward left. The thick, gray colored pulse in each frame indicates the early stage of the perturbation induced by the moving debris which has later transformed. It is important to note that all these evolutions are drawn at the same time evolution (column wise) across different $v_{d}$. The parameters are slightly different for fKdV and mFCT solutions which are optimized respectively to represent this effect. This parameter difference of fKdV and mFCT solutions can be attributed to the weak nonlinearity approximation that is inherent in the KdV formalism and full hydrodynamic simulation through FCT simulation.
@@
We should also note that all these nonlinearities change their behaviors as the plasma evolves in time. So, for example, a steepening event of the ledge edge will eventually transform itself into another form as we evolve the system in time. This is also evident from the similar experimental observations [4, 5]. In this context we would like to recall that a pure KdV equation is integrable with ‘${\rm sech}$’ solution which does not get evolved in time. However, an fKdV equation is inherently non-integrable and different nonlinear solutions of the fKdV equation such as DSW or pinned solitons are necessarily transient phenomena and are bound to evolve in time. As per the parameters used in the FCT simulation, assuming an equilibrium plasma number density of $\sim 10^{6}\,{\rm cm}^{-3}$, the entire spatial nonlinear evolution frame shown in Fig.6 is contained within $\sim 25\,{\rm{\rm cm}}$, which is quite comparable to distance over which these phenomena are observed in contemporary plasma devices [4, 5].
@@ VI Summary and conclusion
To summarize, we have presented a detailed study of external debris-induced nonlinear structures in a multicomponent positive ion-negative ion (PINI) plasma. The motivation to choose a PINI plasma for this study is primarily due the fact that PINI plasmas are easily produced in low-temperature laboratory devices and also are widely used in plasma processing experiments. On the other hand, theoretical and experimental (laboratory and also space-based) studies of debris-induced nonlinear plasma phenomena have been recently generating a lot of interests among the plasma physicists, especially due to the renewed attention of the scientific community in detection and removal of space-debris from low-earth orbits. Besides a detailed analysis of nonlinear waves excited by external charged debris, we also point out the similarities in certain characteristics with those excited by a high velocity positive ion beam or large amplitude perturbations in such plasmas. The important findings of our analysis are presented below.
Figure 6: The changing nonlinearity with increasing debris velocity (from top to bottom) as obtained from the fKdV solutions (left column). The shaded region in each panel indicates the extent of the debris centered at $\xi=0$. The thick, gray pulse in each frame is the debris-induced perturbation at the early stage of the evolution which has later transformed into their respective form as time progresses. The mFCT simulation results are in the right column. Density ratio $(\delta)$ and ion to electron temperature ratio $(\sigma)$ are taken to be $1.05$ and $0.05$ respectively.
@@
Can other nonlinear phenomena mimic debris-induced nonlinearity?
@@
  • Further the amplitudes of nonlinear waves are observed to increase for both positive and negative perturbation with increase in the debris velocity, which is similar to the case of increasing beam velocity in PINI plasma [4, 5].
@@ Acknowledgements.
One of the authors, HS thanks CSIR-HRDG, New Delhi, India for Senior Research Fellowship (SRF) research grant 09/059(0074)/2021-EMR-I.

### v3

@@ V Debris velocity and nonlinearity
One very interesting phenomenon is the changing property of the nonlinearity with increasing debris velocity which is similar to what is observed experimentally with respect to beam velocity in a PINI plasma [12]. From our analysis, what we observe is that as debris velocity increases, the wave amplitude steadily increases until reaching a maximum and then decreases, which is seen from the steepening in the leading edge (precursor region), which then gradually changes to a DSW before becoming a train of solitons and then a single KdV-like soliton. While the steepening in the leading edge indicates a dominant positive nonlinearity over dispersion at low debris velocity, as debris velocity increases, nonlinearity decreases and gets balanced by dispersion which creates a KdV-like soliton.
Figure 6: The changing nonlinearity with increasing debris velocity (from top to bottom) as obtained from the fKdV solutions (left column) and FCT simulation (right column). The thick, gray pulse in each frame is the debris-induced perturbation at the initial stage of the evolution which has later transformed into their respective form as time progresses.
In Fig.6, we show this effect as obtained from the fKdV solution [left column Fig.6(a-d)] and from FCT hydrodynamic simulation [right column Fig.6(e-h)]. In this figure, we have indicated the changing nonlinearity with an arrow with the keywords ‘steepening’, ‘DSW’, ‘train’, and ‘KdV’, respectively to indicate the phenomenon of steepening in the leading edge, formation of DSW, transformation to a train of solitons, and finally a simple KdV soliton as debris velocity $v_{d}$ increases from $0.8$ to $1.5$ for fKdV solutions and from $0.3$ to $1.8$ for FCT solutions. All the panels in the above figure are drawn at the rest frame of the debris, which is moving toward left. The thick, gray colored pulse in each frame indicates the initial stage of the perturbation induced by the moving debris which has later transformed. Density ratio $(\delta)$ and ion to electron temperature ratio $(\sigma)$ are taken to be $1.05$ and $0.05$ respectively, in both cases. It is important to note that all these evolutions are drawn at the same time evolution (column wise) across different $v_{d}$. The parameters are slightly different for fKdV and FCT solutions which are optimized respectively to represent this effect. This parameter difference of fKdV and FCT solutions can be attributed to the weak nonlinearity approximation that is inherent in the KdV formalism and full hydrodynamic simulation through FCT simulation.
@@
We should also note that all these nonlinearities change their behaviors as the plasma evolves in time. For example, a steepening of the leading edge will eventually transform itself into another form as we evolve the system in time. This is also evident from the similar experimental observations [12, 13]. In this context we would like to recall that a pure KdV equation is integrable with ‘${\rm sech}$’ solution which does not get evolved in time. However, an fKdV equation is inherently non-integrable and different nonlinear solutions of the fKdV equation such as DSW or pinned solitons are necessarily transient phenomena and are bound to evolve in time. As per the parameters used in the FCT simulation, assuming an equilibrium plasma number density of $\sim 10^{6}\,{\rm cm}^{-3}$, the entire spatial nonlinear evolution frame shown in Fig.6 is contained within $\sim 25\,{\rm{\rm cm}}$, which is quite comparable to distance over which these phenomena are observed in contemporary plasma devices [12, 13].
@@ VI Summary and conclusion
To summarize, we have presented a detailed study of external debris-induced nonlinear structures in a multicomponent positive ion-negative ion (PINI) plasma. The motivation to choose a PINI plasma for this study is primarily due the fact that PINI plasmas are easily produced in low-temperature laboratory devices and also are widely used in plasma processing experiments. Additionally, theoretical and experimental (laboratory and also space-based) studies of debris-induced nonlinear plasma phenomena have been recently generating a lot of interests among the plasma physicists, especially due to the renewed attention of the scientific community in detection and removal of space-debris from low-earth orbits. Besides a detailed analysis of nonlinear waves excited by external charged debris, we also point out the similarities in certain characteristics with those excited by a high velocity positive ion beam and large amplitude pulse perturbations in such plasmas. The important findings of our analysis are presented below.
@@
Can other processes mimic debris-induced nonlinearity?
@@
  • Further the amplitudes of nonlinear waves are observed to increase for both positive and negative perturbation with increase in the debris velocity, which is similar to the case of increasing beam velocity in PINI plasma [11, 12, 13].
This work advances the understanding of externally driven nonlinear wave dynamics in multicomponent plasmas, offering a broader perspective on the applicability and limitations of perturbative approaches. The observed similarities between debris-driven nonlinear structures and and those due to other processes provide insights into the underlying mechanisms and may help in the interpretation of future experimental results involving negative ion plasmas.
@@ Acknowledgements.
The authors would like to thank Kalpana Bora for critical reading of the revised manuscript and the anonymous referees for their suggestions. One of the authors, HS thanks CSIR-HRDG, New Delhi, India for Senior Research Fellowship (SRF) research grant 09/059(0074)/2021-EMR-I.
Appendix A
In this Appendix, we try to estimate the Landau damping rate for a simplified electron-heavy ion plasma where we have neglected the negative ion species. For this case, using a reductive perturbation theory along with the kinetic correction due to linear Landau damping of the IAW, we have a KdV-like equation [31]
$\partial_{\tau}\phi^{(1)}+\frac{\beta}{\alpha}\phi^{(1)}\partial_{\xi}\phi^{(1)}+\frac{1}{\alpha}\partial_{\xi}^{3}\phi^{(1)}+\tilde{\gamma}\cdot{\cal P}\int_{-\infty}^{+\infty}\frac{\partial\phi^{(1)}}{\partial\xi^{\prime}}\frac{d\xi^{\prime}}{\xi-\xi^{\prime}}=0,$ (A1)
where
$\lambda$ $=$ $1+\frac{3}{2}\sigma,$ (A2)
$\alpha$ $=$ $2\lambda^{-2}-3\lambda^{-4}\sigma,$ (A3)
$\tilde{\gamma}$ $=$ $\frac{\lambda}{\alpha\varepsilon\sqrt{2\pi}}\left[\left(\frac{m_{e}}{m_{+}}\right)^{1/2}+\sigma^{-3/2}\exp\left(-\frac{\lambda^{2}}{2\sigma}\right)\right],$ (A4)
and ${\cal P}$ denotes the principal value integration. The last term with the principal value integration in Eq.(A1) contains the linear Landau damping term of associated IAW and the Eq.(A1) as a whole accounts for weakly nonlinear IAW with Landau damping. We note that integration in Eq.(A1) is in the form of a convolution integral (with of course the principal value) and can be written as
${\cal P}\int_{-\infty}^{+\infty}\left(\frac{\partial\phi}{\partial\xi^{\prime}}\right)\frac{d\xi^{\prime}}{\xi-\xi^{\prime}}={\cal P}\left(\frac{1}{\xi}\right)*\left(\frac{\partial\phi}{\partial\xi^{\prime}}\right),$ (A5)
where ${\cal P}(\xi^{-1})$ is the principal value distribution of $\xi^{-1}$ and the star ‘*’ denotes the convolution operation. Taking Fourier transform ${\cal F}$ of the above expression and applying convolution theorem, we have
${\cal F}\left[{\cal P}\int_{-\infty}^{+\infty}\left(\frac{\partial\phi}{\partial\xi^{\prime}}\right)\frac{d\xi^{\prime}}{\xi-\xi^{\prime}}\right]={\cal F}\left[{\cal P}\left(\frac{1}{\xi}\right)\right]{\cal F}\left[\left(\frac{\partial\phi}{\partial\xi^{\prime}}\right)\right].$ (A6)
The Fourier transform of the terms on the right hand side can be readily evaluated,
${\cal F}\left[{\cal P}\left(\frac{1}{\xi}\right)\right]$ $=$ $-i\pi\cdot{\rm sgn}(k),$ (A7)
${\cal F}\left[\left(\frac{\partial\phi}{\partial\xi^{\prime}}\right)\right]$ $=$ $ik\,\hat{\phi}(k),$ (A8)
where $\hat{\phi}(k)$ is Fourier transform of the function $\phi(\xi)$.
Considering now only the contribution of the Landau damping term, neglecting the nonlinear and the dispersion terms of Eq.(A1), and taking Fourier transform in space and time domain, we obtain the linear dispersion relation with only the Landau damping term as
$\omega=-i\cdot\tilde{\gamma}\cdot\pi k\cdot{\rm sgn}(k)\equiv-i\pi\tilde{\gamma}|k|,$ (A9)
with $(\pi\tilde{\gamma}|k|)$ as the damping rate. Note that the expression for damping rate can also be derived intuitively by comparing the dimensions with that of the dissipation term in a KdV-Burgers equation in a viscous plasma with viscosity coefficient $\nu$
$\left[\tilde{\gamma}\cdot{\cal P}\int_{-\infty}^{+\infty}\frac{\partial\phi^{(1)}}{\partial\xi^{\prime}}\frac{d\xi^{\prime}}{\xi-\xi^{\prime}}\right]=\left[\nu\frac{\partial^{2}\phi^{(1)}}{\partial\xi^{2}}\right],$ (A10)
where the term on the right hand side is the viscous damping term in a KdV-Burgers equation. Note that Eq.(A1) is same as the KdV-Burgers equation with last term replaced by the dissipative damping term. The viscous damping rate of a fluid modeled by the KdV-Burgers equation is given by $(\nu k^{2})$, with $k$ being the wave number of the associated wave, which is in this case is IAW. So, by dimensional analysis, we have
$[\nu k^{2}]\simeq[\tilde{\gamma}k].$ (A11)